# Are LLMs Naturally Good at Synthetic Tabular Data Generation?

### v2

@@ 3.2 Discovery and Distillation of Functional Dependencies (FD)
We leverage FD discovery techniques to govern the order of the autoregressive data generation process in PAFT . A large body of research from the database literature on FD discovery [26, 44, 27] can be leveraged in PAFT , including methods that account for noisy FDs [44]. In this work, we focus on leveraging schema-level FDs discovered using a state-of-the-art FD discovery algorithm [27] to govern the autoregressive data generation process.
@@
We first construct a dependency graph model $G(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ represents the set of vertices with each $v\in\mathcal{V}$ representing an attribute in $\mathcal{S}$ and $e_{ij}\in\mathcal{E}$ representing an edge relation from attribute $v_{i}$ to attribute $v_{j}$. Further, we consider two types of edges in $\mathcal{E}$, specifically each $e_{ij}$ may be a type-1 edge or a type-2 edge (defined next). The two edge types (i.e., type-1, type-2) in $\mathcal{E}$ are derived from three classes of FDs, as shown in Fig. 4(a). Subsequently, we proceed to examine each individual case: 1) Single attribute left-hand side (LHS) and single attribute right-hand side (RHS) 2) Multi-attribute LHS and single-attribute RHS. 3) single or multi-attribute LHS and multi-attribute RHS. We shall use the example table in Fig.  6 to define each FD case.
FD Case 3 relationships are ones where the RHS has multiple attributes and the LHS could have single or multiple-attributes. Such relationships do not directly result in an edge in our dependency graph $\mathcal{G}$. Instead, as classically done in FD literature [27], we subject such FD relationships to an intermediate decomposition step. Specifically, the multi-attribute RHS of FD Case 3 relationships is decomposed into multiple single-attribute RHS dependencies each comprising the original LHS. Further, in each of these new decomposed relationships, if the LHS is single-attribute, it is treated as a FD Case 1 relationship (i.e., a directed edge from $LHS$ to $RHS$ is added to $\mathcal{G}$), else it is handled as an FD Case 2 relationship, wherein for each attribute in the multi-attribute LHS, a backward dependency edge from the single-attribute RHS is added to the dependency graph $\mathcal{G}$. Therefore, for every column in the right-hand side (RHS), we employ either type-1 or type-2 edge construction, depending on the value of its left-hand side (LHS). Full procedure detailed in Appendix A.2 Algorithm 1 .
3.3 Putting It All Together
We define this task of obtaining a total feature order $\mathbf{k}$ from $\mathcal{G}(\mathcal{V},\mathcal{E})$ as an optimization step which seeks to produce $\mathbf{k}$ while minimizing the number of violated relationships in $\mathcal{G}(\mathcal{V},\mathcal{E})$. Appendix A.2 Algorithm 2 outlines this procedure.
@@ 3.3 Putting It All Together
Figure 4: PAFT: FD-Distillation and Depdendency Graph Sorting
3.4 Synthetic Data Generation using PAFT
@@
After the optimized feature order permutation is obtained (via Appendix A.2 Algorithm 2), we fine-tune the LLM with the textually encoded table record $\mathbf{t}_{i}$ such that the auto-regressive generation process is governed by the optimal feature order permutation $\mathbf{k}=\phi(\mathcal{D},S)$. Specifically, we generate the table governed by order $\mathbf{k}$ as defined in Eq. 4.
$\mathbb{P}(\textbf{t}_{i})=\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{j}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}})$ (4)
@@
We employ the Low-Rank Adaptation (LoRA) fine-tuning strategy [16]. To generate synthetic rows, we first sample the initial token $p(t_{i,k_{1}})$ from the marginal distribution of variable $k_{1}$ in actual training data, and then use Eq. 4 to sequentially sample subsequent tokens $p(t_{i,k_{j}})$, where $j\in{2,...,m}$.
4 Experimental Evaluation
In this section, we empirically evaluate PAFT using multiple qualitative and quantitative measures. Specifically, we investigate 7 research questions (see Table 1) to inspect the generation prowess of PAFT with respect to state-of-the-art models while also investigating the degree to which PAFT generated synthetic data is faithful to real-world data distributions and functional dependencies.
ID Research Question Section
RQ1 How do complex data characteristics affect generation quality? Sec. 4.2
RQ2 Does PAFT generate data consistent with intrinsic data characteristics? Sec. 4.3
RQ3 Can data generated by PAFT replace real data in downstream ML model training? Sec. 4.4
RQ4 Does PAFT adhere to real distributions and exhibit mode diversity? Sec. B.1
RQ5 Does the synthetic data generated by PAFT pass the sniff test? Sec. B.2
RQ6 Can PAFT enhance the stability in generating high quality tables, resulting in a faster sampling phase? Sec. B.3
Table 1: Research questions investigated in our work.
4.1 Dataset and Baseline
Dataset. We evaluated the efficiency of PAFT through experiments on six real datasets commonly used in synthetic table generation studies like GReaT [3] CTGAN [40]. These are: Beijing [7], US-locations [13], California Housing [24], Adult Income [2], Seattle [34], and Travel [37]. Separately we also generate a set of four simulated datasets (as described in Algorithm 3 Appendix C.1). Details of dataset statistics are in the Appendix C.1.
Baseline. In benchmarking suite, we have baselines that consist of current deep learning approaches for synthetic data generation (CTGAN [40], CopulaGAN [40], TabSyn [43]), and the most advanced LLM fine-tuning synthetic table generator GReaT [3]. To guarantee an equitable comparison, we employ the Distill-GReaT model for both LLM techniques in all tests. Details in Appendix C.2.
4.2 RQ1: How do complex data characteristics affect generation quality?
Figure 1 has shown that generating a mixture-class of data is more challenging than generating single-class data for baseline methods. However, our approach is capable of effectively handling this task. In order to emphasize the conditional distribution between two columns, we will also assess the performance of our technique and the baseline method for various classes within a mixture-class dataset.
Figure 5: Different states depict different complexity levels for an LLM to fine-tune with. The semantic complexity level increases from left to right. We demonstrate that, in comparison to baselines, PAFT effectively decreases instances of violations (represented by the color orange) in the generated data. Violation rate refers to occurrence of instances that transgress boundaries of a state.
Figure 5 and 10 illustrates that various classes within a complicated dataset may present distinct challenges when it comes to modeling, respectively in terms of visualization and statistics. Both baseline and PAFT can provide satisfactory performance in less complex classes. As the level of complexity rises, the importance of finding the most optimal permutation becomes crucial, and hence we find that PAFT , which is aware of the functional dependencies and hence able to better capture semantic structure in the data, outperforms the GReaT baseline which is agnostic of functional dependencies. The variation may arise due to statistical factors or semantic factors in the actual world, which are examined in a detailed case study (Appendix B.4).
@@ 4.2 RQ1: How do complex data characteristics affect generation quality?
4.3 RQ2: Does PAFT generate data respecting the consistency of intrinsic data characteristics?
We may infer the degree of dependency retention of data’s intrinsic characteristics as another lens through which to evaluate the quality of the data generated by PAFT and baselines. In line with this, we conducted rule checks that were derived from real-world scenarios, and subsequently evaluated the generated data from all DL and LLM baselines, and our PAFT model. Table 2 displays the rate of violation in the generated data. From the table, we observe that PAFT adheres to data’s characteristics more faithfully (i.e., significantly fewer rule violations) than baseline methods, by learning together with functional dependencies.
Dataset Intrinsic Fact CTGAN CopulaGAN TabSyn GReaT PAFT
US-locations State_code $\rightarrow$ bird 94.66$\pm$0.3295.66$\pm$0.170.10$\pm$0.040.30$\pm$0.000.00$\pm$0.00%
US-locations Lat-long $\rightarrow$ State 99.22$\pm$0.0898.51$\pm$0.0721.50$\pm$0.328.16$\pm$0.142.93$\pm$0.15%
California Lat-long $\rightarrow$ CA 47.56$\pm$6.3799.93$\pm$0.008.83$\pm$0.135.42$\pm$0.161.26$\pm$0.06%
@@ 4.3 RQ2: Does PAFT generate data respecting the consistency of intrinsic data characteristics?
Adult education $\rightarrow$ education-num 83.94$\pm$ 1.93% 19.09$\pm$ 0.58% 1.43$\pm$0.041.24$\pm$0.090.46$\pm$ 0.03%
Seattle Zipcode $\rightarrow$ Seattle 0.00$\pm$0.0099.88$\pm$0.000.00$\pm$0.000.00$\pm$0.000.00$\pm$0.00%
Table 2: Datasets exhibit a range of intrinsic characteristics: functional dependencies, range restrictions, and other types of domain knowledge. Cells denote percentage of violations in generated data.
Method Orig. CTGAN CopulaGAN TabSyn GReaT PAFT
Beijing (*) ($\downarrow$) RF 0.41% 2.49$\pm$0.572.15$\pm$0.290.7$\pm$0.010.57$\pm$0.000.52$\pm$0.00%
LR 1.37% 2.23$\pm$0.541.55$\pm$0.211.25$\pm$0.00.97$\pm$0.011.34$\pm$0.01%
NN 0.99% 2.44$\pm$0.742.83$\pm$1.181.01$\pm$0.141.16$\pm$0.560.95$\pm$0.13%
California (*) ($\downarrow$) RF 0.18% 0.65$\pm$0.090.39$\pm$0.010.22$\pm$0.00.25$\pm$0.000.20$\pm$0.00%
LR 0.30% 0.54$\pm$0.10.5$\pm$0.010.30$\pm$0.00.29$\pm$0.000.31$\pm$0.00%
NN 0.34% 0.53$\pm$0.110.47$\pm$0.020.29$\pm$0.020.3$\pm$0.010.27$\pm$0.00%
Seattle (*) ($\downarrow$) RF 0.33% 0.76$\pm$0.340.38$\pm$0.070.30$\pm$0.060.35$\pm$0.010.28$\pm$0.03%
LR 0.29% 0.74$\pm$0.350.32$\pm$0.030.23$\pm$0.040.33$\pm$0.000.29$\pm$0.03%
NN 0.28% 0.71$\pm$0.330.38$\pm$0.080.28$\pm$0.010.33$\pm$0.000.27$\pm$0.01%
US-locations ($\uparrow$) RF 99.95% 7.17$\pm$1.5845.33$\pm$2.8299.99$\pm$0.0199.84$\pm$0.0799.91$\pm$0.03%
LR 46.1% 5.11$\pm$3.1431.08$\pm$1.9243.69$\pm$1.945.65$\pm$0.8649.41$\pm$1.57%
NN 99.85% 7.56$\pm$4.6153.34$\pm$1.5999.64$\pm$0.1798.94$\pm$1.1699.44$\pm$0.28%
Adult ($\uparrow$) RF 84.97% 71.15$\pm$5.5981.33$\pm$1.5383.69$\pm$0.2883.89$\pm$0.4283.06$\pm$0.35%
LR 78.53% 75.68$\pm$0.1478.18$\pm$ 1.53% 78.38$\pm$0.1176.1$\pm$0.2977.24$\pm$0.09%
NN 76.9% 75.69$\pm$0.1076.6$\pm$1.2678.36$\pm$1.2378.23$\pm$1.3179.16$\pm$0.15%
Travel ($\uparrow$) RF 88.95% 56.35$\pm$2.6467.18$\pm$3.084.09$\pm$1.1879.78$\pm$1.2985.19$\pm$1.99%
LR 82.87% 70.17$\pm$17.4679.56$\pm$0.0083.31$\pm$0.2278.34$\pm$2.0282.76$\pm$1.01%
NN 81.77% 71.05$\pm$17.8579.56$\pm$0.0081.88$\pm$1.3380.77$\pm$1.3383.2$\pm$0.9%
@@
Table 3: MLE Performance (%): Comparison of original data to synthetic data. For datasets denoted as (*), we use a regression model for prediction, and calculate MAPE as performance (where lower scores are ideal); For other datasets, classification models are used for prediction and we calculate the accuracy as performance. The best results are marked in bold and the second-best results are underlined. RF: random forests; LR: linear regression: NN: a traditional multi-layer perceptron.
@@
4.4 RQ3: Can data generated by PAFT replace real data in downstream ML model training?

### v3

@@ 4.2. Discovery and Distillation of Functional Dependencies (FD)
We leverage FD discovery techniques to govern the order of the autoregressive data generation process in PAFT . A large body of research from the database literature on FD discovery (Papenbrock et al., 2015; Zhang et al., 2020; Papenbrock and Naumann, 2016) can be leveraged in PAFT , including methods that account for noisy FDs (Zhang et al., 2020). In this work, we focus on leveraging schema-level FDs discovered using a state-of-the-art FD discovery algorithm (Papenbrock and Naumann, 2016) to govern the autoregressive data generation process.
@@
Figure 2. FD-Distillation and Dependency Graph Sorting for automatically extracting order permutations from tables.
We first construct a dependency graph model $G(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ represents the set of vertices with each $v\in\mathcal{V}$ representing an attribute in $\mathcal{S}$ and $e_{ij}\in\mathcal{E}$ representing an edge relation from attribute $v_{i}$ to attribute $v_{j}$. Further, we consider two types of edges in $\mathcal{E}$, specifically each $e_{ij}$ may be a type-1 edge or a type-2 edge (defined next). The two edge types (i.e., type-1, type-2) in $\mathcal{E}$ are derived from three classes of FDs, as shown in Fig. 2(a). Subsequently, we proceed to examine each individual case: 1) Single attribute left-hand side (LHS) and single attribute right-hand side (RHS) 2) Multi-attribute LHS and single-attribute RHS. 3) single or multi-attribute LHS and multi-attribute RHS. We shall use the example table in Fig.  1 to define each FD case.
FD Case 3 relationships are ones where the RHS has multiple attributes and the LHS could have single or multiple-attributes. Such relationships do not directly result in an edge in our dependency graph $\mathcal{G}$. Instead, as classically done in FD literature (Papenbrock and Naumann, 2016), we subject such FD relationships to an intermediate decomposition step. Specifically, the multi-attribute RHS of FD Case 3 relationships is decomposed into multiple single-attribute RHS dependencies each comprising the original LHS. Further, in each of these new decomposed relationships, if the LHS is single-attribute, it is treated as a FD Case 1 relationship (i.e., a directed edge from $LHS$ to $RHS$ is added to $\mathcal{G}$), else it is handled as an FD Case 2 relationship, wherein for each attribute in the multi-attribute LHS, a backward dependency edge from the single-attribute RHS is added to the dependency graph $\mathcal{G}$. Therefore, for every column in the right-hand side (RHS), we employ either type-1 or type-2 edge construction, depending on the value of its left-hand side (LHS).
4.3. Putting It All Together
We define this task of obtaining a total feature order $\mathbf{k}$ from $\mathcal{G}(\mathcal{V},\mathcal{E})$ as an optimization step which seeks to produce $\mathbf{k}$ while minimizing the number of violated relationships in $\mathcal{G}(\mathcal{V},\mathcal{E})$.
4.4. Synthetic Data Generation using PAFT
@@
After the optimized feature order permutation is obtained, we fine-tune the LLM with the textually encoded table record $\mathbf{t}_{i}$ such that the auto-regressive generation process is governed by the optimal feature order permutation $\mathbf{k}=\phi(\mathcal{D},S)$. Specifically, we generate the table governed by order $\mathbf{k}$ as defined in Eq. 4.
(4) $\mathbb{P}(\textbf{t}_{i})=\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{j}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}}).$
@@
We employ the Low-Rank Adaptation (LoRA) fine-tuning strategy (Hu et al., 2021). To generate synthetic rows, we first sample the initial token $p(t_{i,k_{1}})$ from the marginal distribution of variable $k_{1}$ in actual training data, and then use Eq. 4 to sequentially sample subsequent tokens $p(t_{i,k_{j}})$, where $j\in{2,...,m}$.
5. Experimental Evaluation
We conduct an exhaustive empirical evaluation of PAFT to assess its ability to reproduce realistic data distributions, superiority over other competing approaches, and most importantly substitutability of data generated by PAFT in the context of a larger ML pipeline. More specifically, the questions we seek to answer are:
  1. (1)
    Does PAFT-generated synthetic data accurately capture conditional distributions within categories? (Sec 5.1)
  2. (2)
@@ 5. Experimental Evaluation
    Does PAFT generate data respect the consistency of intrinsic data characteristics? (Sec 5.2)
  3. (3)
    Does the synthetic data generated by PAFT pass the sniff test? (Sec 5.3)
  4. (4)
    Can data generated by PAFT replace real data in downstream ML model training? (Sec 5.4)
  5. (5)
    Do the data sets generated by PAFT adhere to real distributions and possess mode diversity? (Sec 5.5)
  6. (6)
    Do newer generations of LLMs obviate the need for PAFT ? (Sec 5.6)
Datasets. We evaluate the effectiveness of PAFT through experiments on six real datasets commonly used in synthetic table generation studies such as GReaT (Borisov et al., 2022) CTGAN (Xu et al., 2019). These are Beijing (Chen, 2017), US-locations (Evans-Bye, 2015), California Housing (Nugent, 2017), Adult Income (Barry Becker, 1996), Seattle (samuel Cortinhas, 2022), and Travel (Tejashvi, 2021). Separately, we also generate a set of four simulated datasets for class-mixture distributions.
Baselines. For benchmarking, we organize baselines that utilize current deep learning approaches for synthetic data generation (CTGAN (Xu et al., 2019), CopulaGAN (Xu et al., 2019), TabSyn (Zhang et al., 2023b)) and the most advanced synthetic table generator with LLM fine-tuning GReaT (Borisov et al., 2022). To guarantee an equitable comparison, we employ the Distill-GReaT model for both LLM techniques in all tests.
Reproducibility. Each baseline (CTGAN, CopulaGAN, TabSyn, GReaT) adheres to the recommended hyperparameters and utilizes officially released API tools: Synthetic Data Vault (Patki et al., 2016) and GReaT (Borisov et al., 2022). For a fair comparison of GReaT and PAFT, the LoRA fine-tuning parameters are set the same as: Lora attention dimension $r=16$, alpha parameter for Lora scaling $lora\_alpha=32$, the names of the modules to apply the adapter to $target\_modules=c\_attn$, the dropout probability for Lora layers $lora\_dropout=0.05$, $bias=none$.
Parameters for MLE and Discriminator Models. We utilize neural network, linear/logistic regression, and random forest models from the Scikit-Learn package for the ML efficiency and discriminator experiments. Every result is evaluated through the 5-fold cross-validation process.
Low-order statistics (Zhang et al., 2023b) of column-wise data distribution and pair-column correlation are calculated with the SDV library  [^2] https://docs.sdv.dev/sdmetrics/reports/quality-report.
5.1. RQ1: Does PAFT-generated synthetic data accurately capture conditional distributions within categories?
Figure 3. For a composite dataset, US-locations, this comparison examines state-specific violation rates across different synthetic data generation approaches. The error bars represent standard deviation. The states on the x-axis are ordered by decreasing violation rates. PAFT significantly reduces state-specific violations in the composite dataset.
Table 2. Violation rates across different categories in the same dataset highlight the effectiveness of PAFT in addressing conditional distribution challenges in mixed-category datasets. Notably, even though baseline models may perform well in the MLE task or distribution evaluation, they often fail in practical boundary checks, such as functional dependencies (FDs). We can see that states with the lowest violation rates are the easiest to model, i.e. rectangular.
States with the Highest Violation Rates ($\downarrow$) (Sorted by LLM baseline: GReaT)
Category (State) MO AK KY FL WV
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real Data 0% 0% 0% 0% 0%
CTGAN 98.0% 99.3% 97.4% 99.9% 100%
CopulaGAN 99.3% 84.6% 99.3% 97.8% 100%
TabSyn 10.5% 17.2% 22.9% 25.1% 28.1%
GReaT 17.1% 18.4% 20.6% 21.9% 22.3%
PAFT 1.9% 5.1% 3.4% 3.4% 3.4%
States with the Lowest Violation Rates ($\downarrow$) (Sorted by LLM baseline: GReaT)
Category (State) CO KS NM SD UT
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real Data 0% 0% 0% 0% 0%
CTGAN 96.5% 97.0% 99.5% 97.5% 99.3%
CopulaGAN 98.3% 98.8% 98.4% 98.3% 98.4%
TabSyn 11.3% 18.4% 6.2% 17.3% 16.6%
GReaT 0.5% 0.9% 1.2% 1.5% 2.1%
PAFT 0.0% 0.3% 0.9% 0.9% 0.0%
Capturing and generating the diversity in a multi-class setting (composite dataset) has been shown to be a challenging practical problem (Margeloiu et al., [n. d.]). In addition to Table 1, Fig. 3 and Table 2 further illustrate the widespread nature of this issue within the same composite dataset, where the GAN-generated results exhibit an almost 100% violation rate. The term ‘composite dataset’ refers to a dataset in which the distributions across different subcategories show significant variances. The LLM model GReaT, which employs random order permutation, shows a significant improvement in maintaining conditional distributions. Nevertheless, its performance remains inconsistent due to unevenness within categories, resulting in violation rates ranging from 0.5% to 22.3%. In contrast, PAFT consistently controls deviations from the real facts of conditional distributions, maintaining them within a range of 0% to 5%. This reliability holds even in challenging subcategory cases where baseline methods underperform.
5.2. RQ2: Does PAFT generate data respecting the consistency of intrinsic data characteristics?
Table 3. Datasets have intrinsic characteristics like functional dependencies, range restrictions, and other domain knowledge. Results are averaged over five random runs.
Intrinsic Fact (Dataset) Fact Violation Rate ($\downarrow$)
CTGAN Cop.GAN TabSyn GReaT PAFT
@@ 5.2. RQ2: Does PAFT generate data respecting the consistency of intrinsic data characteristics?
Lat-long $\rightarrow$ CA (California) 47.6% 99.9% 8.8% 5.4% 1.3%
[…]
$\rightarrow$ education-num (Adult) 83.9% 19.1% 1.4% 1.2% 0.5%
[…]
In addition to the dissimilar sub-category challenge, we also investigate whether unsatisfactory conditional distributions exist in general across various datasets, and whether the PAFT method can address these issues. In line with this, we conducted rule checks that were derived from real-world scenarios and subsequently evaluated the generated data from all models. Table 3 displays the violation rate in the generated data. From the table, we observe that PAFT adheres to data’s characteristics more faithfully (i.e., significantly fewer rule violations) than baseline methods, by learning together with functional dependencies.
5.3. RQ3: Does the synthetic data generated by PAFT pass the sniff test?
Similarly to the analysis conducted in recent work (Borisov et al., 2022) (GReaT), we employ the random forest (RF) algorithm to train discriminators to distinguish real data (labeled True) and synthetically generated data (labeled False). Subsequently, we test performance on an unseen set (consisting of 50% synthetically generated data and 50% real data). In this experiment, scores represent the percentage of correctly classified entities. In this case, an ideal accuracy score would be close to 50%, which means the discrimniator fails to distinguish between real and synthesized data. The scores are shown in Table 4 and indicate that the data generated by PAFT are most indistinguishable from real data, even by powerful discriminative models.
Table 4. Privacy Discriminator Performance. The scores stand for the accuracy for detecting real or fake data, where the ML models are trained using 50% real data and 50% random data. An ideal accuracy score is 50, indicating the model cannot distinguish between real and synthesized data. The best results are marked in bold, the second-best results are underlined. Results are averaged over five random runs.
Data Sniff Test - ML Discriminator Accuracy (Values closest to 50% are best.)
Method CTGAN Co.GAN TabSyn GReaT PAFT
Beijing 99.16% 98.69% 50.97% 51.1% 50.09%
US-locations 99.94% 97.74% 51.97% 50.47% 50.01%
California 98.35% 86.64% 50.64% 53.74% 49.89%
Adult 94.43% 59.82% 51.64% 51.12% 48.75%
Seattle 87.61% 85.7% 50.12% 68.27% 47.21%
Travel 77.96% 74.14% 50.66% 62.49% 48.18%
@@ 5.3. RQ3: Does the synthetic data generated by PAFT pass the sniff test?
Table 5. MLE performance. MLE performance: For datasets with regression tasks (marked *), performance is evaluated using MAPE (where lower scores are better). For datasets with classification tasks, accuracy is used (where higher scores are better). The best results are marked in bold and the second-best results are underlined. Results are averaged over five random runs
. Regression MAPE ($\downarrow$ Over Different Methods Dataset(*) Orig. CTGAN Co.GAN TabSyn GReaT PAFT Beijing RF 0.41% 2.49% 2.15% 0.7% 0.57% 0.52% LR 1.37% 2.23% 1.55% 1.25% 0.97% 1.34% NN 0.99% 2.44% 2.83% 1.01% 1.16% 0.95% Calif. RF 0.18% 0.65% 0.39% 0.22% 0.25% 0.20% LR 0.30% 0.54% 0.5% 0.30% 0.29% 0.31% NN 0.34% 0.53% 0.47% 0.29% 0.3% 0.27% Seattle RF 0.33% 0.76% 0.38% 0.30% 0.35% 0.28% LR 0.29% 0.74% 0.32% 0.23% 0.33% 0.29% NN 0.28% 0.71% 0.38% 0.28% 0.33% 0.27% Classif. Accuracy ($\uparrow$ Over Different Methods Dataset Orig. CTGAN Co.GAN TabSyn GReaT PAFT US-loc. RF 99.95% 7.17% 45.33% 99.99% 99.84% 99.91% LR 46.1% 5.11% 31.08% 43.69% 45.65% 49.41% NN 99.85% 7.56% 53.34% 99.64% 98.94% 99.44% Adult RF 84.97% 71.15% 81.33% 83.69% 83.89% 83.06% LR 78.53% 75.68% 78.18% 78.38% 76.1% 77.24% NN 76.9% 75.69% 76.6% 78.36% 78.23% 79.16% Travel RF 88.95% 56.35% 67.18% 84.09% 79.78% 85.19% LR 82.87% 70.17% 79.56% 83.31% 78.34% 82.76% NN 81.77% 71.05% 79.56% 81.88% 80.77% 83.20%
@@
5.4. RQ4: Can data generated by PAFT replace real data in downstream ML model training?